OpenStudy (anonymous):

3a^2b^3c^5 divided by 8x^4y^3z ?

OpenStudy (anonymous):

how about using some brackets

OpenStudy (anonymous):

then I will solve it for you

OpenStudy (anonymous):

sorry I mean improper notation bothers me if you dont know how to use it I can show you

OpenStudy (anonymous):

sorry, (3a^2b^3c^5) divided by (8x^4y^3z)

OpenStudy (anonymous):

still wrong but better (3a^(2))(b^(3))(c^(5))/ do the bottom properly now

OpenStudy (anonymous):

actually the brackets you used aren't necessary because (3a^(2))(b^(3))(c^(5)) = a number so yeah. Sorry you need to learn this as it makes it really hard for people to answer your question because right now it looks like it is 

OpenStudy (anonymous):

but yeah to get to the question (3a^(2))(b^(3))(c^(5)) / (8x^(4))(y^(3)((z)) Dont be a jerk and actually use proper notation next time anyways x^(-1) = 1/x but you can't simplify this because the variables are not shared

OpenStudy (anonymous):

a, b and c not being equivilant to x, y, or z

OpenStudy (anonymous):

thank you, and sorry I didn't realize to do proper notation

OpenStudy (anonymous):

its fine sorry for stressing it but it is important to know especially if you want to use https://www.wolframalpha.com/ or a graphing/scientific calculator

OpenStudy (anonymous):

using proper notation you could have just inputted your question in and got an answer :)

OpenStudy (anonymous):

to wolframalpha that is

OpenStudy (anonymous):

that would have been a lot easier! ha thanks (:

OpenStudy (anonymous):

the correct notation would be :) ((3a^(2))(b^(3))(c^(5)))/((8x^(4))(y^(3))(z))